OpenStudy (anonymous):

∫[0,1,(3x^(2)+2)^(2)x,] Evaluate the definite integrals

myininaya (myininaya):

\[\int\limits_{0}^{1}x(3x^2+2)^{2}dx\]

myininaya (myininaya):

try to let u=3x^2+2 then du=6x dx or (du)/6=xdx since x=1, then u=3(1)^2+2=5 since x=0, then u=3(0)^2+2=2 so we have \[\int\limits_{2}^{5}(1/6)*u^{2} du\]

OpenStudy (anonymous):

So I don't need to do anything with the exponent on the outside the parenthesis which was actually sipposed to be a three not a two?